# AD-GS: Object-Aware B-Spline Gaussian Splatting for Self-Supervised Autonomous Driving

### v2

@@ 1 Introduction
To address this challenge, self-supervised models for auto-driving scene rendering have been developed [36, 42, 2, 32, 21, 30, 16, 24, 14]. These models aim to reconstruct both scenes and the motions of the contained objects only using images and LiDAR points, thereby eliminating the reliance on manual 3D annotations. However, their performance may still fall short of expectations. Employing neural networks for motion modeling can be an effective approach but may result in high computational overhead and fail to capture fine-grained local movements [36, 30, 32, 21, 16, 14]. On the other hand, using predefined functions, such as trigonometric functions, to model motion can achieve high rendering speeds [2, 24]. Also, these functions can represent the general motions under the noisy self-supervision due to their global fitting properties which optimize all parameters for each training frame. However, they might still have difficulties in handling local details. For scene modeling, auto-driving scenes can be categorized into various components with distinct properties, such as houses, cars, and trees. Previous research has approached scene segmentation for individual modeling using methods like unsupervised learning [30], feature-based self-supervision [36, 32, 24], or fine-grained semantic self-supervision [16, 42, 14, 21]. While these techniques are effective in scene decomposition, the use of the noisy pseudo ground truth can result in reconstruction artifacts.
@@
In this paper, we introduce AD-GS, a novel self-supervised model designed for high-quality auto-driving scene rendering based on Gaussian splatting [9]. To achieve both global and local motion fitting with high precision, we incorporate learnable B-spline curves and B-spline quaternion curves, combined with trigonometric functions, into the deformation representation of dynamic Gaussians. Unlike methods using only trigonometric functions, B-spline curves fit local details by only optimizing the relevant control points instead of all for the corresponding training frames. Additionally, AD-GS simplifies the scene modeling process by segmenting auto-driving scenes into two parts: objects and background, based on the simplified pseudo 2D segmentation. For the object component, we represent objects using dynamic Gaussians and employ bidirectional temporal visibility masks to handle their sudden appearance or disappearance. For the background component, we keep the Gaussians stationary. This decomposition allows AD-GS to reconstruct each component more accurately, resulting in higher rendering quality under noisy pseudo ground truth conditions. To ensure consistent deformation among Gaussians representing the same object, we introduce visibility and physically rigid regularization, which helps to prevent chaotic behavior.
@@
We compare AD-GS with the state-of-the-art self-supervised auto-driving scene rendering models, and our experiments demonstrate the significant performance improvements achieved by AD-GS. Overall the contributions of this paper can be summarized as follows:
@@ 3.1 Overview
We develop AD-GS using Gaussian splatting [9], a point-based rendering model primarily designed for object-level static scenes, offering high rendering quality. The model represents a scene through multiple 3D Gaussians, where each Gaussian is defined by $G=\{\boldsymbol{\mu},S,R,\sigma,\mathbf{c}\}$, representing position, scaling, rotation, opacity, and color, respectively. The color is expressed using spherical harmonic coefficients (SH). These 3D Gaussians can be rendered and trained through differentiable rasterization [44, 39] with ground-truth images. The Gaussians are first projected onto the camera planes using a view transform matrix $W$ and a projective Jacobian matrix $J$.
$\Sigma^{\prime}=JW\Sigma J^{T}W^{T},\Sigma=RSS^{T}R^{T}$ (1)
Finally, a pixel $\mathbf{u}$ on the camera planes can be rendered by $N$ overlapped Gaussians using the $\alpha$-blending as the following formula,
@@
$\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\mathbf{u}-\boldsymbol{\mu}^{p}_{i})^{T}\Sigma^{\prime-1}_{i}(\mathbf{u}-\boldsymbol{\mu}^{p}_{i})}$ (2)
@@
To adapt Gaussian splatting for self-supervised auto-driving scene rendering, we divide all the $N$ Gaussians into two categories: objects $\boldsymbol{\Omega}_{obj}$ and background $\boldsymbol{\Omega}_{bkg},|\boldsymbol{\Omega}_{obj}|+|\boldsymbol{\Omega}_{bkg}|=N$. For the object Gaussians $G\in\boldsymbol{\Omega}_{obj}$, we model their motions by deforming their positions $\boldsymbol{\mu}$ and rotations $R$ over time using B-splines and trigonometric functions. Additionally, we incorporate the bidirectional temporal visibility mask for these Gaussians to handle the sudden appearance and disappearance of the objects. In contrast, the background Gaussians $G\in\boldsymbol{\Omega}_{bkg}$ remain stationary. Notably, the colors of both object and background Gaussians may vary over time. To account for this, similar to previous work [35], we deform the diffuse SH features $\mathbf{z}$ of all Gaussians $G\in\boldsymbol{\Omega}_{obj}\cup\boldsymbol{\Omega}_{bkg}$ using trigonometric functions.
$\mathbf{z}^{\prime}=\mathbf{z}+\sum_{l=1}^{K}\mathbf{f}_{l}\sin(l\pi\cdot t)+\mathbf{h}_{l}\cos(l\pi\cdot t)$ (3)
$K$ is the maximum level of frequency, $\mathbf{f}_{l},\mathbf{h}_{l}$ are the learnable parameters, and $t$ is the timestamp. Furthermore, we represent the distant parts of a scene, such as sky, using a learnable spherical environment map. Our model is optimized in a self-supervised manner with regularization techniques to maintain stability and prevent chaotic behavior.
@@ 3.2 Learnable B-Spline Motion Curve
Learnable B-Spline Curve. The B-spline curve is well known for its ability to provide local control. Given $n+1$ control points $\mathbf{p}_{i},i\in\{0,1,...,n\}$, we can construct a $C^{k-1}$-continuous $k$-th order ($k\leq n+1$) B-spline curve using the following formula
@@
where $t_{0},t_{1},...,t_{n+1}$ are the knots arranged in non-decreasing order. In practice, we define $t$ as the timestamp, with $t_{k-1}$ and $t_{n+1}$ representing the minimum and maximum timestamps of the input sequence, respectively. The basis function $B_{i,k}$ only has a non-zero value in $[t_{i},t_{i+k}]$. Furthermore, all control points $\mathbf{p}_{i}$ are treated as learnable parameters, and we utilize uniform B-spline curves for motion modeling. However, the de Boor-Cox recursion formula for the basis function $B_{i,k}$ is computationally inefficient. Therefore, we use the matrix formulation of uniform B-spline curves for better computational complexity [25].
@@
$u=\frac{t-t_{i}}{t_{i+1}-t_{i}},u\in[0,1),i=k-1,k,...,n$ (5)
$M_{k}$ is a $k\times k$ matrix that can be precomputed before training, and the explicit expression can be found in Section A of our supplementary. Using Equation 3.2, we can obtain the matrix representation of the B-spline curve for each segment $[t_{i},t_{i+1})$ in Equation 4.
@@
$u=\frac{t-t_{i}}{t_{i+1}-t_{i}},u\in[0,1),i=k-1,k,...,n$ (6)
Unlike MLPs and trigonometric functions, which optimize all parameters for every output, B-spline curves exhibit local control. Each position on the B-spline curve is only influenced by the nearby control points, as illustrated in Figure 2. This property allows the B-spline curve to effectively capture motions with local details by only fine-tuning the relevant control points. However, while the B-spline curve excels at modeling local details, it has limitations in capturing global motion, which is also crucial for accurate reconstruction under the noisy self-supervision. To address this, we combine trigonometric functions with B-splines to model the movement of each object Gaussian $G\in\boldsymbol{\Omega}_{obj}$, achieving both global and local fitting.
$\boldsymbol{\mu}^{\prime}=\boldsymbol{\mu}+\mathbf{p}(t)+\sum_{l=1}^{L}\mathbf{a}_{l}\sin(t\cdot l\pi)+\mathbf{b}_{l}\cos(t\cdot l\pi)$ (7)
$L$ is the maximum frequency level, and $\mathbf{a}_{l},\mathbf{b}_{l}$ are the learnable parameters.
Learnable B-Spline Quaternion Curve. Because the interpolation of unit quaternions is uneven, the standard B-spline curve cannot adequately represent the unit quaternion deformation of Gaussian rotation parameters. To address this limitation, we adopt the B-spline quaternion curve [11] to model Gaussian rotations. Similar to Equation 4, given $n+1$ learnable unit quaternions $\mathbf{q}_{i},i\in\{0,1,...,n\}$ as control points, the $k$-th order unit quaternion B-spline curve in each segment is defined as,
@@
$t\in[t_{i},t_{i+1}),i=k-1,k,...,n$ (8)
where $\mathbf{w}_{i}=\log(\mathbf{q}_{i-1}^{-1}\mathbf{q}_{i})$, and $\widetilde{B}_{i,k}(t)=\sum_{j=i}^{n}B_{j,k}(t)$, which can be efficiently computed using Equation 3.2. Notably, the operators in Equation 3.2 are specifically defined for unit quaternions. Although the B-spline quaternion curve has limitations in global control, we find that directly modeling the rotation of each object Gaussian $G\in\boldsymbol{\Omega}_{obj}$ by setting $R=\mathbf{q}(t)$ is sufficient for our task.
@@ 3.3 Object-Aware Splatting with Temporal Mask
Scene Decomposition with Simplified 2D Segmentation. Previous works employ instance segmentation to guide scene decomposition [16, 21], but this approach might be excessively noisy. To address this, we simplify the segmentation classes into two categories: objects and background. Classes likely to exhibit motion over time, such as cars, are categorized as objects, while all other classes are treated as background. This simplification is sufficient and significantly more robust for rendering scenes in autonomous driving applications. We initialize the Gaussians using LiDAR points and partition them into two subsets $\boldsymbol{\Omega}_{obj}$ and $\boldsymbol{\Omega}_{bkg}$, based on the 2D projected positions on the simplified binary segmentation results $\mathcal{M}_{obj}$ from SAM [12, 26]. Using this initial decomposition, we can render the object mask through the $\alpha$-blending by the following formula:
^$\hat{\mathcal{M}}_{obj}=\sum_{i=1}^{N}\mathbb{I}\{G_{i}\in\boldsymbol{\Omega}_{obj}\}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (9)
@@
$\sigma^{\prime}(t)=\sigma\cdot e^{-\frac{(t-\mu_{t})^{2}}{2s^{2}}},s=\left\{\begin{array}[]{cc}s_{0},&t<\mu_{t}\\ s_{1},&t\geq\mu_{t}\end{array}\right.$ (12)
where $\mu_{t}$ is the fixed acquisition timestamp of the LiDAR points, $s_{0},s_{1}$ are the learnable scales of the bidirectional visibility mask. Unlike previous approaches [2], which treat $\mu_{t}$ as a learnable parameter, we argue that the acquisition timestamp itself serves as an effective prompt to determine the visible moment of an object and distinguish the different objects that have passed through the same place. To prevent the temporal mask from becoming excessively narrow, we design an expanding regularization loss as follows:
$\mathcal{L}_{s}=||\frac{2\Delta_{f}}{s_{0}+s_{1}}||_{1}$ (13)
@@ 3.4 Self-Supervision with Regularization
$\hat{X}_{r}=\sum_{i=1}^{N}\boldsymbol{\mu}_{i}^{\prime}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (14)
where $\boldsymbol{\mu}^{\prime}$ is the position of each Gaussian at the target moment. We then project the $\hat{X}_{r}$ onto the target image, and supervise it using the corresponding pseudo flow points $X_{t}$ through $\mathcal{L}_{f}=||\phi_{t}(\hat{X}_{r})-X_{t}||_{1}$, where $\phi_{t}$ is the projection function. To enhance efficiency and reduce noise, we only supervise the object part of the flow with the simplified 2D pseudo segmentation results obtained from SAM [26, 12].
Inverse Depth Supervision. In prior work, the depth map is rendered through the $\alpha$-blending of the distances between the camera and Gaussian centers, which is then inverted to supervise against the monocular pseudo ground truth. However, when no Gaussian exists along a pixel ray, the inversion operation can become computationally unstable. To address this, we directly render the expectation map of inverse depth [10] by using the following formula:
$\frac{1}{\hat{d}}=\sum_{i=1}^{N}\frac{1}{d_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (15)
where $d_{i}$ is the distance between camera and each Gaussian center. Then we supervise the rendered inverse depth map with the monocular pseudo ground truth $\frac{1}{d}$ generated by DPTv2 [37], using the scale-and-shift depth loss [40].
$\mathcal{L}_{d}=||w\frac{1}{\hat{d}}+q-\frac{1}{d}||_{1}$ (16)
$w,q$ are the optimal values that minimize $argmin_{w,q}(w\frac{1}{\hat{d}}+q-\frac{1}{d})^{2}$. This depth loss incorporates 3D information from sparse viewpoints, enabling AD-GS to reconstruct large auto-driving scenes with greater precision.
Visibility and Physically Rigid Regularization. To satisfy the assumption of local rigidity and prevent AD-GS from becoming disorganized, we introduce a regularization loss to ensure that nearby object Gaussians $G\in\boldsymbol{\Omega}_{obj}$ exhibit similar deformations in position and temporal visibility mask across the timeline.
$\mathcal{L}_{r}=\sum_{\boldsymbol{\beta}\in\{\mathbf{p},\mathbf{a},\mathbf{b},\{s_{0},s_{1}\}\}}\sum_{\beta_{i}\in\boldsymbol{\beta}}var(\beta_{i})$ (17)
$var$ means the parameter variance of the 8 nearby Gaussians selected by the KNN algorithm. To address the high computational cost of the KNN operation, we cache the KNN results and update them every 10 iterations, striking a balance between efficiency and performance.
@@
$\lambda_{r}\mathcal{L}_{r}+\lambda_{s}\mathcal{L}_{s}$ (18)
where $\lambda$ are the weight hyperparameters for each loss, $\mathcal{L}_{1},\mathcal{L}_{D-SSIM}$ are the supervision using the ground truth images [9], and $\mathcal{L}_{sky}=BCE(1-O,\mathcal{M}_{sky})$ is the binary cross-entropy loss between the accumulated opacity $O=\prod_{i=1}^{N}(1-\alpha_{i})$ and the sky mask $\mathcal{M}_{sky}$ generated by SAM [12, 26]. The pipeline of AD-GS is shown in Figure 3.
Table 1: Quantitative comparisons on the KITTI [7] dataset. We follow the experimental setup of SUDS [32], and the color of each cell shows the best and the second best. ’Annotations’ means whether the model is assisted by the manual 3D annotations.
KITTI-$75\%$ KITTI-$50\%$ KITTI-$25\%$
@@
Table 2: Quantitative comparisons on the Waymo [29] dataset. We mainly follow the experimental setup of StreetGS [35] with a higher resolution $1280\times 1920$, and the color of each cell shows the best and the second best. ’Annotations’ means whether the model is assisted by the manual 3D annotations. denotes the metric only for moving objects.

### v3

@@ 1 Introduction
To address this challenge, self-supervised models for auto-driving scene rendering have been developed [36, 42, 2, 32, 21, 30, 16, 24, 14]. These models aim to reconstruct both scenes and motions of the contained objects only using images and LiDAR points, thereby eliminating the reliance on manual 3D annotations. However, their performance may still fall short of expectations. Employing neural networks for motion modeling can be an effective approach but may result in high computational overhead and fail to capture the fine-grained local movements of objects [36, 30, 32, 21, 16, 14]. On the other hand, using predefined functions, such as trigonometric functions, to model motion can achieve high rendering speeds [2, 24]. Also, these functions can represent the general motions under the noisy self-supervision due to their global fitting properties which optimize all parameters for each training frame. However, they might still have difficulties in handling the local details of motions. For scene modeling, auto-driving scenes can be categorized into various components with distinct properties, such as houses, cars, and trees. Previous research has approached scene segmentation for independent modeling using methods like unsupervised learning [30], feature-based self-supervision [36, 32, 24], or fine-grained semantic self-supervision [16, 42, 14, 21]. While these techniques are effective in scene decomposition, the use of the noisy pseudo ground truth can result in reconstruction artifacts.
@@
In this paper, we introduce AD-GS, a novel self-supervised model designed for high-quality auto-driving scene rendering based on Gaussian splatting [9]. To achieve both global and local motion fitting with high precision, we incorporate learnable B-spline curves and B-spline quaternion curves, combined with trigonometric functions, into the deformation representation of dynamic Gaussians. Unlike the methods using only trigonometric functions [2, 24], B-spline curves fit local details by only optimizing the relevant control points instead of all for the corresponding training frames. Additionally, AD-GS simplifies the scene modeling process by segmenting auto-driving scenes into two parts: objects and background, based on the simplified pseudo 2D segmentation. For the object component, we represent objects using dynamic Gaussians and employ bidirectional temporal visibility masks to handle their sudden appearance or disappearance. For the background component, we keep the Gaussians stationary. This decomposition allows AD-GS to reconstruct each component more accurately, resulting in higher rendering quality under noisy pseudo ground truth conditions. To ensure consistent deformation among Gaussians representing the same object, we introduce visibility and physically rigid regularization, which helps to prevent chaotic behavior.
@@
We compare AD-GS with the state-of-the-art self-supervised auto-driving scene rendering models, and our experiments demonstrate the significant performance improvements achieved by AD-GS. Overall the contributions of this paper can be summarized as follows,
@@ 3.1 Overview
We develop AD-GS using Gaussian splatting [9], a point-based rendering model primarily designed for object-level static scenes, offering high rendering quality. The model represents a scene through multiple 3D Gaussians, where each Gaussian is defined by $G=\{\boldsymbol{\mu},S,R,\sigma,\mathbf{c}\}$, representing position, scaling, rotation, opacity, and color, respectively. The color is expressed using spherical harmonic coefficients (SH). These 3D Gaussians can be rendered and trained through differentiable rasterization [44, 39] with ground-truth images. The Gaussians are first projected onto the camera planes using a view transform matrix $W$ and a projective Jacobian matrix $J$:
$\Sigma^{\prime}=JW\Sigma J^{T}W^{T},\Sigma=RSS^{T}R^{T}.$ (1)
Finally, a pixel $\mathbf{u}$ on the camera planes can be rendered by $N$ overlapped Gaussians using the $\alpha$-blending as the following formula:
@@
$\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\mathbf{u}-\boldsymbol{\mu}^{p}_{i})^{T}\Sigma^{\prime-1}_{i}(\mathbf{u}-\boldsymbol{\mu}^{p}_{i})},$ (2)
@@
To adapt Gaussian splatting for self-supervised auto-driving scene rendering, we divide all the $N$ Gaussians into two categories: objects $\boldsymbol{\Omega}_{obj}$ and background $\boldsymbol{\Omega}_{bkg},|\boldsymbol{\Omega}_{obj}|+|\boldsymbol{\Omega}_{bkg}|=N$. For the object Gaussians $G\in\boldsymbol{\Omega}_{obj}$, we model their motions by deforming their positions $\boldsymbol{\mu}$ and rotations $R$ over time using B-splines and trigonometric functions. Additionally, we incorporate the bidirectional temporal visibility mask for these Gaussians to handle the sudden appearance and disappearance of the objects. In contrast, the background Gaussians $G\in\boldsymbol{\Omega}_{bkg}$ remain stationary. Notably, the colors of both object and background Gaussians may vary over time. To account for this, similar to previous work [35], we deform the diffuse SH features $\mathbf{z}$ of all Gaussians $G\in\boldsymbol{\Omega}_{obj}\cup\boldsymbol{\Omega}_{bkg}$ using trigonometric functions:
$\mathbf{z}^{\prime}=\mathbf{z}+\sum_{l=1}^{K}\mathbf{f}_{l}\sin(l\pi\cdot t)+\mathbf{h}_{l}\cos(l\pi\cdot t).$ (3)
$K$ is the maximum level of frequency, $\mathbf{f}_{l}$ and $\mathbf{h}_{l}$ are the learnable parameters, and $t$ is the timestamp. Furthermore, we represent the distant parts of a scene, such as sky, using a learnable spherical environment map. Our model is optimized in a self-supervised manner with regularization techniques to maintain stability and prevent chaotic behavior. The pipeline of AD-GS is shown in Figure 3.
@@ 3.2 Learnable B-Spline Motion Curve
Learnable B-Spline Curve. The B-spline curve is well known for its ability to provide local control. Given $n+1$ control points $\mathbf{p}_{i},i\in\{0,1,...,n\}$, we can construct a $C^{k-1}$-continuous $k$-th order ($k\leq n+1$) B-spline curve using the following formula,
@@
where $t_{0},t_{1},...,t_{n+1}$ are the knots arranged in non-decreasing order. In practice, we define $t$ as the timestamp, with $t_{k-1}$ and $t_{n+1}$ representing the minimum and maximum timestamps of the input sequence, respectively. The basis function $B_{i,k}$ only has a non-zero value in $t_{i},t_{i+k}]$. Furthermore, all control points $\mathbf{p}_{i}$ are treated as learnable parameters, and we utilize uniform B-spline curves for motion modeling. However, the de Boor-Cox recursion formula for the basis function $B_{i,k}$ is computationally inefficient. Therefore, we use the matrix formulation of uniform B-spline curves for better computational complexity [25],
@@
$u=\frac{t-t_{i}}{t_{i+1}-t_{i}},u\in[0,1),i=k-1,k,...,n.$ (5)
$M_{k}$ is a $k\times k$ matrix that can be precomputed before training, and the explicit expression can be found in Section A of our supplementary. Using Equation 3.2, we can obtain the matrix representation of the B-spline curve for each segment $t_{i},t_{i+1})$ in Equation 4:
@@
$u=\frac{t-t_{i}}{t_{i+1}-t_{i}},u\in[0,1),i=k-1,k,...,n.$ (6)
Unlike MLPs and trigonometric functions, which optimize all parameters for every output, B-spline curves exhibit local fitting. Each position on the B-spline curve is only influenced by the nearby control points, as illustrated in Figure 2. This property allows the B-spline curve to effectively capture motions with local details by only fine-tuning the relevant control points. However, while the B-spline curve excels at modeling local details, it has limitations in capturing global motion, which is also crucial for accurate reconstruction under the noisy self-supervision. To address this, we combine trigonometric functions with B-splines to model the movement of each object Gaussian $G\in\boldsymbol{\Omega}_{obj}$, achieving both global and local fitting:
$\boldsymbol{\mu}^{\prime}=\boldsymbol{\mu}+\mathbf{p}(t)+\sum_{l=1}^{L}\mathbf{a}_{l}\sin(t\cdot l\pi)+\mathbf{b}_{l}\cos(t\cdot l\pi),$ (7)
where $L$ is the maximum frequency level, $\mathbf{a}_{l}$ and $\mathbf{b}_{l}$ are the learnable parameters.
Learnable B-Spline Quaternion Curve. Because the interpolation of unit quaternions is uneven, the standard B-spline curve cannot adequately represent the unit quaternion deformation of Gaussian rotation parameters. To address this limitation, we adopt the B-spline quaternion curve [11] to model Gaussian rotations. Similar to Equation 4, given $n+1$ learnable unit quaternions $\mathbf{q}_{i},i\in\{0,1,...,n\},$ as control points, the $k$-th order unit quaternion B-spline curve in each segment is defined as,
@@
$t\in[t_{i},t_{i+1}),i=k-1,k,...,n,$ (8)
where $\mathbf{w}_{i}=\log(\mathbf{q}_{i-1}^{-1}\mathbf{q}_{i})$, and $\widetilde{B}_{i,k}(t)=\sum_{j=i}^{n}B_{j,k}(t)$, which can be efficiently computed using Equation 3.2. Notably, the operators in Equation 3.2 are specifically defined for unit quaternions. Although the B-spline quaternion curve has limitations in global fitting, we find that directly modeling the rotation of each object Gaussian $G\in\boldsymbol{\Omega}_{obj}$ by setting $R=\mathbf{q}(t)$ is sufficient for our task.
@@ 3.3 Object-Aware Splatting with Temporal Mask
Scene Decomposition with Simplified 2D Segmentation. Previous works employ instance segmentation to guide scene decomposition [16, 21], but this approach might perform excessively noisy. Therefore, we simplify the segmentation classes into two categories: objects and background. Classes likely to exhibit motion over time, such as cars, are categorized as objects, while all other classes are treated as background. This simplification is sufficient and significantly more robust for rendering scenes in autonomous driving applications. We initialize the Gaussians using LiDAR points and partition them into two subsets $\boldsymbol{\Omega}_{obj}$ and $\boldsymbol{\Omega}_{bkg}$, based on the 2D projected positions on the simplified binary segmentation results $\mathcal{M}_{obj}$ from SAM [12, 26]. Using this initial decomposition, we can render the object mask through the $\alpha$-blending by the following formula:
$\hat{\mathcal{M}}_{obj}=\sum_{i=1}^{N}\mathbb{I}\{G_{i}\in\boldsymbol{\Omega}_{obj}\}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (9)
@@
$\sigma^{\prime}(t)=\sigma\cdot e^{-\frac{(t-\mu_{t})^{2}}{2s^{2}}},s=\left\{\begin{array}[]{cc}s_{0},&t<\mu_{t}\\ s_{1},&t\geq\mu_{t}\end{array}\right$., (12)
where $\mu_{t}$ is the fixed acquisition timestamp of the LiDAR points, $s_{0},s_{1}$ are the learnable scales of the bidirectional visibility mask. Unlike previous approaches [2], which treat $\mu_{t}$ as a learnable parameter, we argue that the acquisition timestamp itself serves as an effective prompt to determine the visible moment of an object and distinguish the different objects that have passed through the same location. To prevent the temporal mask from becoming excessively narrow, we design an expanding regularization loss as follows:
[…]
@@ 3.4 Self-Supervision with Regularization
$\hat{X}_{r}=\sum_{i=1}^{N}\boldsymbol{\mu}_{i}^{\prime}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (14)
[…]
Inverse Depth Supervision. In prior work, the depth map is rendered through the $\alpha$-blending of the distances between the camera and Gaussian centers, which is then inverted to supervise against the monocular pseudo ground truth. However, when no Gaussian exists along a pixel ray, the inversion operation can become computationally unstable. To address this problem, we directly render the expectation map of inverse depth [10] by using the following formula:
$\frac{1}{\hat{d}}=\sum_{i=1}^{N}\frac{1}{d_{i}}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (15)
where $d_{i}$ is the distance between camera and each Gaussian center. Then we supervise the rendered inverse depth map with the monocular pseudo ground truth $\frac{1}{d}$ generated by DPTv2 [37], using the scale-and-shift depth loss [40]:
[…]
$w,q$ are the optimal values that minimize $(w\frac{1}{\hat{d}}+q-\frac{1}{d})^{2}$. This depth loss incorporates 3D information from sparse viewpoints, enabling AD-GS to reconstruct large auto-driving scenes more accurately.
Visibility and Physically Rigid Regularization. To satisfy the assumption of local rigidity and prevent AD-GS from becoming disorganized, we introduce a regularization loss to ensure that nearby object Gaussians $G\in\boldsymbol{\Omega}_{obj}$ exhibit similar deformations in position and temporal visibility mask across the timeline:
$\mathcal{L}_{r}=\sum_{\boldsymbol{\beta}\in\{\mathbf{p},\mathbf{a},\mathbf{b},\{s_{0},s_{1}\}\}}\sum_{\beta_{i}\in\boldsymbol{\beta}}var(\beta_{i}).$ (17)
varvar means the parameter variance of the nearby 8 Gaussians selected by the KNN algorithm. To address the high computational cost of the KNN operation, we cache the KNN results and update them every 10 iterations, striking a balance between efficiency and performance.
@@
$\lambda_{r}\mathcal{L}_{r}+\lambda_{s}\mathcal{L}_{s},$ (18)
where $\lambda$ are the weight hyperparameters for each loss, $\mathcal{L}_{1},\mathcal{L}_{D-SSIM}$ are the supervision using the ground truth images [9], and $\mathcal{L}_{sky}=BCE(1-O,\mathcal{M}_{sky})$ is the binary cross-entropy loss between the accumulated opacity $O=\prod_{i=1}^{N}(1-\alpha_{i})$ and the sky mask $\mathcal{M}_{sky}$ generated by SAM [12, 26].
Table 1: Quantitative comparisons on the KITTI [7] dataset. We follow the experimental setup of SUDS [32], and the color of each cell shows the best and the second best. “Annotations” means whether the model is assisted by the manual 3D annotations.
$75\% KITTI-$$50\%$ KITTI-$25\%$
@@
Table 2: Quantitative comparisons on the Waymo [29] dataset. We mainly follow the experimental setup of StreetGS [35] with a higher resolution $1280\times 1920$, and the color of each cell shows the best and the second best. “Annotations” means whether the model is assisted by the manual 3D annotations. denotes the metric only for moving objects.